Gucchi:

The angles shown below are supplementary: What is the value of x? 10 degrees 11 degrees 20 degrees 21 degrees

Gucchi:

1 attachment
Gucchi:

@florisalreadytaken

NewtIXC:

try this equation: 70+10x=180 solve for x

Florisalreadytaken:

\[ 10x+70= 180^\circ \] \[ 10x=110^\circ \] \[ x= \frac{110}{10} \] \[ \Huge x=?? \]

Gucchi:

@florisalreadytaken wrote:
\[ 10x+70= 180^\circ \] \[ 10x=110^\circ \] \[ x= \frac{110}{10} \] \[ \Huge x=?? \]
11

Florisalreadytaken:

there you go!

Gucchi:

thank you!